OpenStudy (anonymous):

Simplify sqrt(27) - sqrt(3) + 2[sqrt(12)]

OpenStudy (anonymous):

= 3sqrt(3) - sqrt(3) + 4sqrt(3) =6sqrt(3)

OpenStudy (anonymous):

√(27) - √(3) + 2√(12) √(9)(3) - √(3) + 2√(4)(3) 3√3 - 1√3 + 4√3 2√3 + 4√3 6√3

OpenStudy (anonymous):

Would that also mean 6 times the square root of 3

OpenStudy (anonymous):

\[\sqrt{3*9}-\sqrt{3}+2\sqrt{4*3}\] \[3\sqrt{3}-\sqrt{3}+4\sqrt{3}\] \[6\sqrt{3}\]

OpenStudy (anonymous):

@yokie: yes